OpenStudy (anonymous):

Solve the following system of equations and explain which method you chose to solve the system and why you felt it was the best choice. x - y = 13 2x + y = 2

OpenStudy (mertsj):

Do you want to use substitution or elimination?

OpenStudy (anonymous):

Substitution

OpenStudy (mertsj):

Solve the first equation for x

OpenStudy (anonymous):

How do I do that?

OpenStudy (mertsj):

Add y to both sides of the equation

OpenStudy (anonymous):

Like this? x(y) - y = 13(y) 2x(y) + y = 2(y)

OpenStudy (mertsj):

What is the first equation?

OpenStudy (anonymous):

x - y = 13

OpenStudy (mertsj):